OpenStudy (anonymous):

Which expression is equivalent to 3(c + b) + 9b – c 3c + 9b 2c + 12b 2c -12b 3c – 9b

OpenStudy (id21):

2c+12b

OpenStudy (id21):

3c+3b+9b-c=2c+12b

OpenStudy (anonymous):

\[3\left( c +b\right) +9b -c = 2c + 12b\]

OpenStudy (anonymous):

2c + 12b